. Kelly's art teacher is modeling a project for art and algebra.The students are going to create a project with blocks in a patteren. Students will use algebra to describe the pattern. A student puts two formulas on the board. Recursive:f(0)=3, f(n)=2*f(n-1)
Explict: f(n)=3*2^n
a. using the recursive formula above,write the first 5 terms of the sequence.
b.You want to know the value of f(30). Is it more efficent to use the recursive or explicit formula? Explain your answer.

Answer by stanbon(75887)

A student puts two formulas on the board. Recursive:f(0)=3, f(n)=2*f(n-1)
Explict: f(n)=3*2^n
a. using the recursive formula above,write the first 5 terms of the sequence.
f(0) = 3
f(1) = 2*f(0) = 2*3 = 6
f(2) = 2*f(1) = 2*2*3 = 2^2*3 = 12
f(3) = 2^3*3 = 8*3 = 24
f(4) = 2^4*3 = 48
================================
b.You want to know the value of f(30).
f(30) = 2^30*3 = 3221225472
-------------------------------------
Is it more efficent to use the recursive or explicit formula? Explain your answer.
I'll leave that to you.
Cheers,
Stan H.
==================